# Fedor V. Fomin

### v2

@@ 15                         $O\leftarrow$ empty graph over the vertex set $\tau(T_{i})$;
23                         $M_{i}\leftarrow O$; $M\leftarrow M$ where $M[T_{i}]$ is replaced by $M_{i}$;
@@ 33            $\mathcal{M}\leftarrow\mathcal{M}\cup\{(M,f)\}$;
Finally, we estimate the running time. First note that all lines except Algorithm 1 and Algorithm 1 incur at most a polynomial factor. By Lemma 4, up to polynomial factors Algorithm 1 takes time $2^{(h+s+k)^{\mathcal{O}(k)}}$, for each invocation. This holds since $|V(H_{i})|+|E(H_{i})|\leq 3h+3s$, see the proof of 14. Moving to Algorithm 1, let $w=h+2s$ be the maximum size of $W$; there are at most $w^{t}$ choices for $\xi$, and at most $|V(H)|^{s+w}$ choices for $f_{W}$. We may assume $t\leq k$, since otherwise, an independent set of size $k$ can be found trivially by taking an arbitrary vertex from each $C_{i}$. We now argue that with a fixed $f_{W}$, there are at most $(s+w)^{2(s+w)}$ choices for $M_{W}$ that satisfy CD1. Indeed, terminals are fixed by $f_{W}$, therefore, it only remains to set the edges where at least one endpoint is a non-terminal. Since every non-terminal has degree two in $M_{W}$, there are at most $(s+w)^{2}$ choices for each non-terminal, and at most $(s+w)^{2(s+w)}$ choices in total. This part can also be upper-bounded by $(s+h)^{\mathcal{O}(s+h)}=2^{(s+h)^{1+o(1)}}$. The overall running time is therefore at most $2^{(s+h+k)^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}$. ∎
@@ 16            $S\leftarrow S\cup S^{\prime}$;
We iteratively construct the vertex set $S$ suitable for Lemma 5, then invoke Lemma 5. We start by setting $S=\emptyset$, then iteratively find a separator in one of the connected components of $G-S$, and add it to $S$. This happens until the connected components of $G-S$ are all $\kappa$-connected, where $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$. If, at some point, there are at least $k$ connected components, we return the independent set constructed by taking a vertex from each connected component. Otherwise, when the connectivity condition is fulfilled, we run Lemma 5 on the given input with the computed $S$, and return the output. See Algorithm 2 for a detailed algorithm description.
We now argue that the algorithm is correct. If the algorithm returns at Algorithm 2, then the output is an independent set in $G$ of size $k$. Otherwise, the algorithm eventually reaches Algorithm 2, since at every iteration of the main loop, the size of $S$ increases by at least one. It is left to verify that the call to Lemma 5 is valid. For that, we need that every connected component of $G-S$ is $\kappa$-connected, for $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$. However, this is exactly the condition at Algorithm 2, and this is the only option for the algorithm to break out of the main loop while reaching Algorithm 2. Therefore, the algorithm’s output is correct by Lemma 5.
It remains to upper-bound the running time of the algorithm. For that, it is crucial to upper-bound the size of $S$ when the algorithm reaches Algorithm 2. First, we observe that the main loop performs at most $k$ iterations before it stops either at Algorithm 2 or at Algorithm 2. Specifically, at the start of the $j$-th iteration, the number of connected components in $G-S$ is at least $j$, which implies that at iteration $k$, the loop necessarily stops at Algorithm 2. This holds since each iteration increases the number of connected components by at least one, since a separator of one of the former components is added into $S$, while the remaining components are unchanged.
@@
Since the number of iterations is at most $k$, the size of $S$ by Algorithm 2 is then at most $s_{k}=k^{\mathcal{O}(k)}\cdot h$. We now upper-bound the running time of Algorithm 1 by the running time bound of Lemma 5, since clearly up to Algorithm 2 the algorithm takes polynomial time. By Lemma 5, the total running time is thus at most $2^{(hk^{\mathcal{O}(k)})^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}=2^{|H|^{\mathcal{O}(k)}}\cdot 2^{k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$. ∎
5 Conclusion
In conclusion, this paper shows that Hamiltonian Path/Cycle and Path Cover are $\operatorname{{\sf FPT}}$ parameterized by $\alpha(G)$. Furthermore, Path Cover is $\operatorname{{\sf FPT}}$ parameterized “below” the Gallai–Milgram bound.
We conclude with several open research questions. All our results concern undirected graphs. The complexity of all these problems on directed graphs of small independence number remains open. For example, we do not know whether Hamiltonian Cycle is $\operatorname{{\sf NP}}$-complete on directed graphs with an independence number $k$ for some fixed $k\geq 2$. But we cannot exclude that the problem may be in $\operatorname{{\sf X}P}$ or even $\operatorname{{\sf FPT}}$ parameterized by $k$. Parameterized complexity offers a powerful toolbox for obtaining lower bounds for structural parameterization [CNP+11, FGL+19, FGLS10, LMS18]. However, all these reductions use gadgets with large independence numbers. Obtaining algorithmic lower bounds for parameterization with small independence number appears more challenging.

### v3

@@ 15          $O\leftarrow$ empty graph over the vertex set $\tau(T_{i})$;
23          $_{i}\leftarrow O; MM$$\leftarrow M$ where $M[T_{i}]$ is replaced by $M_{i}$;
@@ 33    $\mathcal{M}\leftarrow\mathcal{M}\cup\{(M,f)\}$;
Finally, we estimate the running time. First note that all lines except Algorithm 1 and Algorithm 1 incur at most a polynomial factor. By Lemma 4, up to polynomial factors Algorithm 1 takes time $2^{(h+s+k)^{\mathcal{O}(k)}}$, for each invocation. This holds since $|V(H_{i})|+|E(H_{i})|\leq 3h+3s$, see the proof of 14. Moving to Algorithm 1, let $w=h+2s$ be the maximum size of $W$; there are at most $t^{w}$ choices for $\xi$, and at most ${(s+w)}^{|V(H)|}$ choices for $f_{W}$. We may assume $t\leq k$, since otherwise, an independent set of size $k$ can be found trivially by taking an arbitrary vertex from each $C_{i}$. We now argue that with a fixed $f_{W}$, there are at most $(s+w)^{2(s+w)}$ choices for $M_{W}$ that satisfy CD1. Indeed, terminals are fixed by $f_{W}$, therefore, it only remains to set the edges where at least one endpoint is a non-terminal. Since every non-terminal has degree two in $M_{W}$, there are at most $(s+w)^{2}$ choices for each non-terminal, and at most $(s+w)^{2(s+w)}$ choices in total. In total, Algorithm 1 performs at most $(s+h+k)^{\mathcal{O}(s+h)}=2^{(s+h+k)^{1+o(1)}}$ iterations. The overall running time is therefore at most $2^{(s+h+k)^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}$. ∎
@@ 16    $S\leftarrow S\cup S^{\prime}$;
We iteratively construct the vertex set $S$ suitable for Lemma 5, then invoke Lemma 5. We start by setting $S=\emptyset$, then iteratively find a separator in one of the connected components of GSG-S, and add it to $S$. This happens until the connected components of GSG-S are all $\kappa$-connected, where $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$. If, at some point, there are at least $k$ connected components, we return the independent set constructed by taking a vertex from each connected component. Otherwise, when the connectivity condition is fulfilled, we run Lemma 5 on the given input with the computed $S$, and return the output. See Algorithm 2 for a detailed algorithm description.
We now argue that the algorithm is correct. If the algorithm returns at Algorithm 2, then the output is an independent set in $G$ of size $k$. Otherwise, the algorithm eventually reaches Algorithm 2, since at every iteration of the main loop, the size of $S$ increases by at least one. It is left to verify that the call to Lemma 5 is valid. For that, we need that every connected component of GSG-S is $\kappa$-connected, for $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$. However, this is exactly the condition at Algorithm 2, and this is the only option for the algorithm to break out of the main loop while reaching Algorithm 2. Therefore, the algorithm’s output is correct by Lemma 5.
It remains to upper-bound the running time of the algorithm. For that, it is crucial to upper-bound the size of $S$ when the algorithm reaches Algorithm 2. First, we observe that the main loop performs at most $k$ iterations before it stops either at Algorithm 2 or at Algorithm 2. Specifically, at the start of the $j$-th iteration, the number of connected components in GSG-S is at least $j$, which implies that at iteration $k$, the loop necessarily stops at Algorithm 2. This holds since each iteration increases the number of connected components by at least one, since a separator of one of the former components is added into $S$, while the remaining components are unchanged.
@@
Since the number of iterations is at most $k$, the size of $S$ by Algorithm 2 is then at most $s_{k}=k^{\mathcal{O}(k)}\cdot h$. We now upper-bound the running time of Algorithm 1 by the running time bound of Lemma 5, since clearly up to Algorithm 2 the algorithm takes polynomial time. By Lemma 5, the total running time is thus at most $2^{(hk^{\mathcal{O}(k)})^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}=2^{|H|^{\mathcal{O}(k)}\cdot k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$. ∎
5 Open questions
We conclude with several open research questions. All our results concern undirected graphs. The complexity of all these problems on directed graphs of small independence number remains open. For example, we do not know whether Hamiltonian Cycle is $\operatorname{{\sf NP}}$-complete on directed graphs with an independence number $k$ for some fixed $k\geq 2$. But we cannot exclude that the problem may be in $\operatorname{{\sf X}P}$ or even $\operatorname{{\sf FPT}}$ parameterized by $k$. Parameterized complexity offers a powerful toolbox for obtaining lower bounds for structural parameterization [23, 36, 35, 73]. However, all these reductions use gadgets with large independence numbers. Obtaining algorithmic lower bounds for parameterization with small independence number appears more challenging.
@@ 5 Open questions
Acknowledgements
Fedor V. Fomin and Petr A. Golovach were supported by the Research Council of Norway via the BWCA project (grant no. 314528), Fedor V. Fomin was supported by the European Research Council (ERC) via grant NewPC (reference 101199930). [Uncaptioned image]
Petr A. Golovach was supported by the Research Council of Norway via the Extreme-Algorithms project (grant no 355137).
Nikola Jedličková and Jan Kratochvíl were supported by GAUK 370122. In the later stages of preparing this article, Nikola Jedličková was supported by the project PRIMUS/24/SCI/008 and Charles University Research Centre program No. UNCE/24/SCI/022.
Danil Sagunov was supported by the Ministry of Science and Higher Education of the Russian Federation (agreement 075-15-2025-344 dated 29/04/2025 for Saint Petersburg Leonhard Euler International Mathematical Institute at PDMI RAS).